Nathan has two brothers, Conner and Josh. If we add 2 to conners age, it is exactly 1/2 the age of their older brother josh who is 12, How old is Conner?
That question then leads into this one: How many years will it be until Conner's age is half of josh's age.

I know the answers to both (4 on both) but can not figure out how to write it into an equation. PLEASE help :)

Answer by drk(1908)

Let N = nathan, C = Conner, J = Josh
C + 2 = (1/2)Josh.
Josh is 12, so
C + 2 = (1/2)*12 = 6
So, Conner is 4.
Conner is 4 and josh is 12.
C + a = (1/2)(J + a)
4 + a = (1/2)(12 + a)
4 + a = 6 + (1/2)a
2 = (3/2)a
4 = a.
In 4 years c = 1/2 josh age